# Weakly-Supervised 3D Visual Grounding based on Visual Linguistic Alignment

### v3

@@ Abstract
Learning to ground natural language queries to target objects or regions in 3D point clouds is quite essential for 3D scene understanding. Nevertheless, existing 3D visual grounding approaches require a substantial number of bounding box annotations for text queries, which is time-consuming and labor-intensive to obtain. In this paper, we propose 3D-VLA, a weakly supervised approach for 3D visual grounding based on Visual Linguistic Alignment. Our 3D-VLA exploits the superior ability of current large-scale vision-language models (VLMs) on aligning the semantics between texts and 2D images, as well as the naturally existing correspondences between 2D images and 3D point clouds, and thus implicitly constructs correspondences between texts and 3D point clouds with no need for fine-grained box annotations in the training procedure. During the inference stage, the learned text-3D correspondence will help us ground the text queries to the 3D target objects even without 2D images. To the best of our knowledge, this is the first work to investigate 3D visual grounding in a weakly supervised manner by involving large scale vision-language models, and extensive experiments on ReferIt3D and ScanRefer datasets demonstrate that our 3D-VLA achieves comparable and even superior results over the fully supervised methods.
@@ I Introduction
To solve the above challenge, a natural way is to investigate 3D visual grounding in a weakly supervised manner that does not need dense object-sentence annotations. Such idea has been explored in 2D visual grounding, which mainly focus on establishing semantic correspondences between 2D image and text descriptions [22, 23, 24, 66]. However, different from 2D images, 3D point clouds inherently provide essential geometric information and surface context with a higher level of complexity and a larger spatial scale, and bring new challenges to effectively learning the matching relationships between 3D point clouds and texts. Wang et al[63] approach weakly supervised 3D visual grounding as a coarse-to-fine semantic matching problem and propose a coarse-to-fine semantic matching model to build the relationship between 3D point cloud and text. But we can find that the relationship is not well-aligned and not as expectation. Therefore how to correlate texts and 3D point clouds is still a big challenge. But we notice that, the correspondences between 3D point clouds and 2D images can be easily obtained by geometric camera calibration with intrinsic and extrinsic parameters. At the same time, we can also note that the current pre-trained large-scale vision-language models (VLMs) such as CLIP [25], ViLT [10], VLMO [11] have been greatly developed. Using massive text-image pairs for model training, VLMs are able to establish precise semantic matching relationships between natural languages and 2D images, and have achieved good results in various downstream tasks such as image classification [52, 53], visual question answering [54], and image captioning [55]. So, as shown in Fig.1 (b), why don’t we take 2D images as a bridge, leveraging the correspondences between point clouds and images, images and natural languages, to implicitly build matching relationships between point clouds and natural languages?
@@
To this end, we present a novel weakly supervised method 3D-VLA, which explores the 3D visual grounding based on the Visual Linguistic Alignment while without the need of 3D bounding box annotations. Specifically, as shown in Fig. 2, in the training stage, our proposed 3D-VLA possesses a text module, a 2D module, and a 3D module. We first extract 3D proposal candidates from the point cloud scene and project these proposals to 2D image regions through geometric camera calibration, and then we utilize a frozen CLIP model to get the embeddings of the text query and 2D image regions with its text encoder and image encoder, respectively. The correspondences between the text query and 2D image regions can thus be measured through their CLIP embeddings. We leverage contrastive learning to optimize the 3D encoder in the 3D module by making the learned 3D embeddings comparable to the text and 2D CLIP embeddings. If a 2D image region and a 3D proposal are matched in pairs, their embeddings should be pulled closer, otherwise they should be pushed further apart.
@@
Ideally, if the 3D embedding of a proposal candidate is learned well enough, it can be directly compared with the text query embedding by the similarity measurement to judge whether it is the target proposal. However, we also find that only relying on the implicit contrastive learning is not reliable since the pretrained data of VLMs are general and do not have specialized knowledge to the indoor point cloud scene. We propose to alleviate this problem by introducing multi-modal adaption through task-aware classification. As shown in Fig. 2, we first add three adapters to transfer the text, 2D and 3D embeddings to another embedding space, and then the 2D and 3D classification are realized by comparing the adapted region/proposal embeddings to the text embeddings of the category labels in the dataset. For the query, we directly apply a text classifier on its adapted query embedding, thus obtaining its distribution on the category labels. By introducing the task-aware classification signal of 3D visual grounding in the indoor point cloud scene, we can further align the semantic relationships among texts, 2D images and 3D point clouds specialized for 3D visual grounding.
@@ II-A Weakly Supervised Visual Grounding on Images
However, owing to proposals generated in the first stage is of low quality, the performances of those models are limited. To address the issue of imprecise object proposals generated in the first stage, some one-stage pipeline [31, 33, 34] are introduced. For example, 3D-SPS [33] directly performs 3D visual grounding at a single stage and treats 3D visual grounding task as a keypoint selection problem to find the most target-related keypoints. In order to well align visual-language feature, Wu et al[34] propose a text decoupling module to parse linguistic description into multiple semantic components.
Those methods mentioned above are all fully supervised, which need much expensive bounding box annotations. To overcome this shortcoming, Wang et al[63] propose a coarse-to-fine semantic matching model to build the relationship of each proposal and text. Owing to lacking of the annotation of 3D bounding boxes which are realted to the text, the relationship it build is not well-aligned and the performance of it is not as expectation. Therefore, how to build the well-aligned relationship between 3D point cloud and text is still a problem.
@@ III Method
In this section, we will first demonstrate our 3D-VLA training procedure by visual linguistic alignment. Then, we will describe the inference procedure of 3D-VLA with category-oriented proposal filtering.
III-A 3D-VLA Training by Visual Linguistic Alignment
@@ III-A2 Text Encoder
For each 3D proposal candidate, we project its point clouds onto $L$ sampled frames [42] in the original video through geometric camera calibration, and get the corresponded 2D image regions. Actually, we find that each 3D proposal may have multiple correspondences in different frames in the video and therefore refer to multiple 2D image regions. Here, we only choose the 2D image region which contains the most 3D projected points from the point cloud, to pair with the 3D proposal candidate. We leverage the image encoder of CLIP $\varepsilon^{2d}$ to extract the 2D semantic embeddings of these 2D image regions, which are denoted as $F^{2D}=\left\{F_{1}^{2D},\dots,F_{M}^{2D}\right\}$. Similarly, we also freeze $\varepsilon^{2d}$ and directly load the CLIP pretrained parameters.
@@ III-A5 Multi-Modal Adaption Through Task-Aware Classification
Fully Supervised ReferIt3D [5] ECCV20 3D 53.75% 37.47% 21.03% 12.83% 26.44% 16.90%
ScanRefer [12] ECCV20 3D 67.64% 46.19% 32.06% 21.26% 38.97% 26.10%
3D+2D 76.33% 53.51% 32.73% 21.11% 41.19% 27.40%
TGNN [19] AAAI21 3D 68.61% 56.80% 29.84% 23.18% 37.37% 29.70%
InstanceRefer [20] ICCV21 3D 77.45% 66.83% 31.27% 24.77% 40.23% 32.93%
SAT [38] ICCV21 3D+2D 73.21% 50.83% 37.64% 25.16% 44.54% 30.14%
3D-SPS [33] CVPR22 3D+2D 84.12% 66.72% 40.32% 29.82% 48.82% 36.98%
EDA [34] CVPR23 3D 85.76% 68.57% 49.13% 37.64% 54.59% 42.26%
HAM [15] - 3D 79.24% 67.86% 41.46% 34.03% 48.79% 40.60%
M3DRef-CLIP [16] ICCV23 3D - 77.2% - 36.8% - 44.7%
ConcreteNet [17] - 3D 82.39% 75.62% 41.24% 36.56% 48.91% 43.84%
Weakly Supervised Wang et al[63] ICCV23 3D - - - - 27.37% 21.96%
Ours - 3D+2D 72.95% 62.17% 22.77% 17.94% 32.51% 26.53%
@@
Fully Supervised ReferIt3D [5] ECCV20 35.6%±0.7% 43.6%±0.8% 27.9%±0.7% 32.5%±0.7% 37.1%±0.8%
TGNN [19] AAAI21 37.3%±0.3% 44.2%±0.4% 30.6%±0.2% 35.8%±0.2% 38.0%±0.3%
InstanceRefer [20] ICCV21 38.8%±0.4% 46.0%±0.5% 31.8%±0.4% 34.5%±0.6% 41.9%±0.4%
SAT [38] ICCV21 49.2%±0.3% 56.3%±0.5% 42.4%±0.4% 46.9%±0.3% 50.4%±0.3%
LanguageRefer [45] CoRL22 43.9% 51.0% 36.6% 41.7% 45.0%
3D-SPS [33] CVPR22 51.5%±0.2% 58.1%±0.3% 45.1%±0.4% 48.0%±0.2% 53.2%±0.3%
BUTD-DETR [31] ECCV22 54.6% 60.7% 48.4% 46.0% 58.0%
EDA [34] CVPR23 52.1% - - - -
HAM [15] - 48.2% 54.3% 41.9% 41.5% 51.4%
Weakly Supervised Ours - 32.1%±0.2% 38.6%±0.2% 25.8%±0.3% 28.8%±0.3% 33.7%±0.4%
@@
Fully Supervised ReferIt3D [5] ECCV20 40.8%±0.2% 44.7%±0.1% 31.5%±0.4% 39.2%±1.0% 40.8%±0.1%
TGNN [19] AAAI21 45.0%±0.2% 48.5%±0.2% 36.9%±0.5% 45.8%±1.1% 45.0%±0.2%
InstanceRefer [20] ICCV21 48.0%±0.3% 51.1%±0.2% 40.5%±0.3% 45.4%±0.9% 48.1%±0.3%
SAT [38] ICCV21 57.9% 61.2% 50.0% 49.2% 58.3%
LanguageRefer [45] CoRL22 56.0% 58.9% 49.3% 49.2% 56.3%
3D-SPS [33] CVPR22 62.6%±0.2% 56.2%±0.6% 65.4%±0.1% 49.2%±0.5% 63.2%±0.2%
BUTD-DETR [31] ECCV22 67.0% 68.6% 63.2% 53.0% 67.6%
EDA [34] CVPR23 68.1% - - - -
HAM [15] - 62.5% 65.9% 54.6% 52.5% 63.0%
Weakly Supervised Ours - 34.5%±0.2% 37.7%±0.2% 27.0%±0.4% 35.3%±0.5% 34.5%±0.2%
@@
Wang et al[63] ICCV23 27.3% 21.1% 18.0% 14.4% 21.6% 16.8% 22.9% 18.1% 22.5% 17.6%
Ours - 33.3% 25.3% 24.2% 17.3% 30.3% 20.8% 31.7% 21.6% 28.7% 21.3%
@@
Wang et al[63] ICCV23 29.4% 24.9% 21.0% 17.5% 20.2% 17.2% 27.2% 22.9% 26.9% 22.7%
Ours - 35.2% 28.1% 25.8% 21.1% 27.3% 22.3% 33.5% 27.4% 30.5% 24.8%
@@
1 31.4%±0.2% 38.9%±0.4% 24.2%±0.5% 28.2%±0.4% 33.0%±0.2%
2 31.8%±0.2% 38.5%±0.4% 25.3%±0.5% 29.7%±0.4% 32.8%±0.4%
3 32.1%±0.2% 38.6%±0.2% 25.8%±0.3% 28.8%±0.3% 33.7%±0.4%
4 31.7%±0.3% 38.9%±0.5% 24.8%±0.1% 28.5%±0.2% 33.3%±0.4%
TABLE VI: Performance comparison with some baselines
@@
Ours 38.6% 25.8% 28.8% 33.7% 32.1%
Random 2.0% 2.0% 1.9% 2.0% 2.0%
Hard Weakly Supervised 35.8% 22.4% 28.4% 29.2% 29.0%
Fully Supervised 43.1% 30.6% 30.9% 39.6% 36.7%

### v4

# Weakly-Supervised 3D Visual Grounding based on Visual Language Alignment

@@ Abstract
Learning to ground natural language queries to target objects or regions in 3D point clouds is quite essential for 3D scene understanding. Nevertheless, existing 3D visual grounding approaches require a substantial number of bounding box annotations for text queries, which is time-consuming and labor-intensive to obtain. In this paper, we propose 3D-VLA, a weakly supervised approach for 3D visual grounding based on Visual Language Alignment. Our 3D-VLA exploits the superior ability of current large-scale vision-language models (VLMs) on aligning the semantics between texts and 2D images, as well as the naturally existing correspondences between 2D images and 3D point clouds, and thus implicitly constructs correspondences between texts and 3D point clouds with no need for fine-grained box annotations in the training procedure. During the inference stage, the learned text-3D correspondence will help us ground the text queries to the 3D target objects even without 2D images. To the best of our knowledge, this is the first work to investigate 3D visual grounding in a weakly supervised manner by involving large scale vision-language models, and extensive experiments on ReferIt3D and ScanRefer datasets demonstrate that our 3D-VLA achieves comparable and even superior results over the fully supervised methods.The code will be available at https://github.com/xuxiaoxxxx/3D-VLA.
@@ I Introduction
To solve the above challenge, a natural way is to investigate 3D visual grounding in a weakly supervised manner that does not need dense object-sentence annotations. Such idea has been explored in 2D visual grounding, which mainly focus on establishing semantic correspondences between 2D image and text descriptions [22, 23, 24, 72]. However, different from 2D images, 3D point clouds inherently provide essential geometric information and surface context with a higher level of complexity and a larger spatial scale, and bring new challenges to effectively learning the matching relationships between 3D point clouds and texts.
Wang et al[63] utilize a coarse-to-fine matching method with contrastive learning to identify top-$k$ candidate proposals, followed by text reconstruction loss for supervision. However, the low quality of candidate proposals selected in the first stage, coupled with reconstruction losses supervised solely by text embeddings, results in a alignment between the 3D and text relationships that is inadequate and not as expectation. Therefore how to correlate texts and 3D point clouds is still a big challenge.
We can also notice that, the correspondences between 3D point clouds and 2D images can be easily obtained by geometric camera calibration with intrinsic and extrinsic parameters. At the same time, we can also note that the current pre-trained large-scale vision-language models (VLMs) such as CLIP [25], ViLT [10], VLMO [11] have been greatly developed. Using massive text-image pairs for model training, VLMs are able to establish precise semantic matching relationships between natural languages and 2D images, and have achieved good results in various downstream tasks such as image classification [52, 53], visual question answering [54], and image captioning [55]. So, as shown in Fig.1 (b), why don’t we take 2D images as a bridge, leveraging the correspondences between point clouds and images, images and natural languages, to implicitly build matching relationships between point clouds and natural languages?
@@
To this end, we present a novel weakly supervised method 3D-VLA, which explores the 3D visual grounding based on the Visual Language Alignment while without the need of 3D bounding box annotations. Specifically, as shown in Fig. 2, in the training stage, our proposed 3D-VLA possesses a text module, a 2D module, and a 3D module. We first extract 3D proposal candidates from the point cloud scene and project these proposals to 2D image regions through geometric camera calibration, and then we utilize a frozen CLIP model to get the embeddings of the text query and 2D image regions with its text encoder and image encoder, respectively. The correspondences between the text query and 2D image regions can thus be measured through their CLIP embeddings. We leverage contrastive learning to optimize the 3D encoder in the 3D module by making the learned 3D embeddings comparable to the text and 2D CLIP embeddings. If a 2D image region and a 3D proposal are matched in pairs, their embeddings should be pulled closer, otherwise they should be pushed further apart.
@@
Ideally, if the 3D embedding of a proposal candidate is learned well enough, it can be directly compared with the text query embedding by the similarity measurement to judge whether it is the target proposal. However, we observe that relying solely on implicit contrastive learning is unreliable, as the pretrained data of VLMs is general and lacks specialized knowledge for indoor point cloud scenes. Indoor environments present greater complexity, characterized by higher object density and intricate spatial relationships, making accurate visual grounding when using only VLMs and contrastive learning methods. Therefore, we propose to alleviate this problem by introducing multi-modal adaption through task-aware classification. As shown in Fig. 2, we first add three adapters to transfer the text, 2D and 3D embeddings to another embedding space, and then the 2D and 3D classification are realized by comparing the adapted region/proposal embeddings to the text embeddings of the category labels in the dataset. For the query, we directly apply a text classifier on its adapted query embedding, thus obtaining its distribution on the category labels. By introducing the task-aware classification signal of 3D visual grounding in the indoor point cloud scene, we can further align the semantic relationships among texts, 2D images and 3D point clouds specialized for 3D visual grounding.
@@ II-A Weakly Supervised Visual Grounding on Images
However, owing to proposals generated in the first stage is of low quality, the performances of those models are limited. To address the issue of imprecise object proposals generated in the first stage, some one-stage pipeline [31, 33, 34] are introduced. For example, 3D-SPS [33] directly performs 3D visual grounding at a single stage and treats 3D visual grounding task as a keypoint selection problem to find the most target-related keypoints. In order to well align visual-language feature, Wu et al[34] propose a text decoupling module to parse language description into multiple semantic components.
Those methods mentioned above are all fully supervised, which need much expensive bounding box annotations. To overcome this shortcoming, Wang et al[63] adopt a two-stage coarse-to-fine semantic matching approach. In the first stage, they use contrastive learning to align 3D object and sentence query features, selecting top-$k$ object candidates based on whether the object-query pairs within the same scene are positive or negative. In the second stage, a semantic reconstruction module is introduced to compute the fine-grained semantic similarity between 3D objects and the sentence query, selecting the target object with the lowest reconstruction loss from the candidates. However, in the one hand, in large 3D scenes with many objects, this two-stage matching process struggles to achieve precise alignment, especially in complex environments where object features overlap. In the other hand, The reconstruction loss is supervised solely by the text embeddings, which are too weak to reliably guide the model in selecting the best-matching object. These reasons cause that the relationship it build is not well-aligned and the performance of it is not as expectation. Therefore, how to build the well-aligned relationship between 3D point cloud and text is still a problem.
@@ III Method
In this section, we will first demonstrate our 3D-VLA training procedure by visual language alignment. Then, we will describe the inference procedure of 3D-VLA with category-oriented proposal filtering.
III-A 3D-VLA Training by Visual Language Alignment
@@ III-A2 Text Encoder
For each 3D proposal candidate, we project its point clouds onto $L$ sampled frames [42] in the original video through geometric camera calibration, and get the corresponded 2D image regions. To avoid the potential inaccuracies in the 2D-3D correspondences, we apply a boundary extension strategy after projecting the 3D point cloud onto 2D space. Specifically, we expand the projected 2D region $[x,y,w,h]$ by 10% along both the width and height, i.e., $[x,y,w+0.2*w,h+0.2*h]$, to account for potential deviations caused by the projection. This strategy helps to capture the correct region more reliably, even when minor projection errors occur. Actually, we find that each 3D proposal may have multiple correspondences in different frames in the video and therefore refer to multiple 2D image regions. Here, we only choose the 2D image region which contains the most 3D projected points from the point cloud, to pair with the 3D proposal candidate. We leverage the image encoder of CLIP $\varepsilon^{2d}$ to extract the 2D semantic embeddings of these 2D image regions, which are denoted as $F^{2D}=\left\{F_{1}^{2D},\dots,F_{M}^{2D}\right\}$. Similarly, we also freeze $\varepsilon^{2d}$ and directly load the CLIP pretrained parameters.
@@ III-A5 Multi-Modal Adaption Through Task-Aware Classification
Fully Supervised ReferIt3D [5] ECCV20 3D 53.75 37.47 21.03 12.83 26.44 16.90
ScanRefer [12] ECCV20 3D 67.64 46.19 32.06 21.26 38.97 26.10
3D+2D 76.33 53.51 32.73 21.11 41.19 27.40
TGNN [19] AAAI21 3D 68.61 56.80 29.84 23.18 37.37 29.70
InstanceRefer [20] ICCV21 3D 77.45 66.83 31.27 24.77 40.23 32.93
SAT [38] ICCV21 3D+2D 73.21 50.83 37.64 25.16 44.54 30.14
3D-SPS [33] CVPR22 3D+2D 84.12 66.72 40.32 29.82 48.82 36.98
EDA [34] CVPR23 3D 85.76 68.57 49.13 37.64 54.59 42.26
HAM [15] - 3D 79.24 67.86 41.46 34.03 48.79 40.60
M3DRef-CLIP [16] ICCV23 3D - 77.2 - 36.8 - 44.7
ConcreteNet [17] - 3D 82.39 75.62 41.24 36.56 48.91 43.84
3D-VisTA [64] ICCV23 3D 77.40 70.90 38.70 34.80 45.90 41.50
G3-LQ [65] CVPR24 3D 88.59 73.28 50.23 39.72 55.95 44.72
Zero Shot LERF [66] ICCV23 3D+2D - - - - 4.4 0.3
Openscene [67] CVPR23 3D+2D - - - - 14.3 4.7
LLM-Grounder [68] ICRA24 3D+2D - - - - 17.1 5.3
Weakly Supervised Wang et al[63] ICCV23 3D - - - - 27.37 21.96
Ours - 3D+2D 72.95 62.17 22.77 17.94 32.51 26.53
@@
Fully Supervised ReferIt3D [5] ECCV20 35.6±0.7 43.6±0.8 27.9±0.7 32.5±0.7 37.1±0.8
TGNN [19] AAAI21 37.3±0.3 44.2±0.4 30.6±0.2 35.8±0.2 38.0±0.3
InstanceRefer [20] ICCV21 38.8±0.4 46.0±0.5 31.8±0.4 34.5±0.6 41.9±0.4
SAT [38] ICCV21 49.2±0.3 56.3±0.5 42.4±0.4 46.9±0.3 50.4±0.3
LanguageRefer [45] CoRL22 43.9 51.0 36.6 41.7 45.0
3D-SPS [33] CVPR22 51.5±0.2 58.1±0.3 45.1±0.4 48.0±0.2 53.2±0.3
BUTD-DETR [31] ECCV22 54.6 60.7 48.4 46.0 58.0
EDA [34] CVPR23 52.1 - - - -
HAM [15] - 48.2 54.3 41.9 41.5 51.4
3D-VisTA [64] ICCV23 57.5 - 49.4 - -
G3-LQ [65] CVPR24 58.4 - 50.7 - -
Weakly Supervised Ours - 32.1±0.2 38.6±0.2 25.8±0.3 28.8±0.3 33.7±0.4
@@
Fully Supervised ReferIt3D [5] ECCV20 40.8±0.2 44.7±0.1 31.5±0.4 39.2±1.0 40.8±0.1
TGNN [19] AAAI21 45.0±0.2 48.5±0.2 36.9±0.5 45.8±1.1 45.0±0.2
InstanceRefer [20] ICCV21 48.0±0.3 51.1±0.2 40.5±0.3 45.4±0.9 48.1±0.3
SAT [38] ICCV21 57.9 61.2 50.0 49.2 58.3
LanguageRefer [45] CoRL22 56.0 58.9 49.3 49.2 56.3
3D-SPS [33] CVPR22 62.6±0.2 56.2±0.6 65.4±0.1 49.2±0.5 63.2±0.2
BUTD-DETR [31] ECCV22 67.0 68.6 63.2 53.0 67.6
EDA [34] CVPR23 68.1 - - - -
HAM [15] - 62.5 65.9 54.6 52.5 63.0
3D-VisTA [64] ICCV23 69.6 - 63.6 - -
G3-LQ [65] CVPR24 73.1 - 66.3 - -
Weakly Supervised Ours - 34.5±0.2 37.7±0.2 27.0±0.4 35.3±0.5 34.5±0.2
@@
Wang et al[63] ICCV23 27.3 21.1 18.0 14.4 21.6 16.8 22.9 18.1 22.5 17.6
Ours - 33.3 25.3 24.2 17.3 30.3 20.8 31.7 21.6 28.7 21.3
@@
Wang et al[63] ICCV23 29.4 24.9 21.0 17.5 20.2 17.2 27.2 22.9 26.9 22.7
Ours - 35.2 28.1 25.8 21.1 27.3 22.3 33.5 27.4 30.5 24.8
@@
1 31.4±0.2 38.9±0.4 24.2±0.5 28.2±0.4 33.0±0.2
2 31.8±0.2 38.5±0.4 25.3±0.5 29.7±0.4 32.8±0.4
3 32.1±0.2 38.6±0.2 25.8±0.3 28.8±0.3 33.7±0.4
4 31.7±0.3 38.9±0.5 24.8±0.1 28.5±0.2 33.3±0.4
TABLE VI: Performance comparison with model variants on Nr3D.
@@
Ours 38.6 25.8 28.8 33.7 32.1
RPS. 2.0 2.0 1.9 2.0 2.0
CBWS. 35.8 22.4 28.4 29.2 29.0
GTS. 43.1 30.6 30.9 39.6 36.7
TABLE VII: Performance comparison with different projection strategy on Nr3D.
Method Overall Easy Hard View-dep. View-indep.
Unmodified Projection 31.1 38.6 23.9 28.9 32.2
Boundary-Extended Projection 32.1 38.6 25.8 28.8 33.7
TABLE VIII: The runtime of three datasets.
NR3D SR3D ScanRefer
Getting 2D image regions (for a room) 336.6s 362.9s 242.3s
Inference (for a query) 0.382s 0.384s 0.397s